OpenStudy (anonymous):

Algebra II: You Sell coffee that coffee A that sells for 11$ a lb. you also have coffee B tht sells for 17$ lb . You want a mixture that is 30lbs that sells for 15$ lb. How much coffee do u need to buy

OpenStudy (anonymous):

Coffee A sells - $11.00 lb Coffe B sells- $17.00 lb You want- 30 lb mixture of both that sells for $15.00lb How much do you want to buy - X

OpenStudy (anonymous):

30 pounds at 15 a pound means the total price will be $450 lets call the amount of coffee A as \(x\) then since the total is 30 the amount of $17 coffee must be \(30-x\) the cost of the $11 coffee will be \(11x\) and the cost of the $17 coffee will bet \(17(30-x)\) since we know the total cost is $450 we can write \[11x+17(30-x)=450\] and solve for \(x\)

OpenStudy (anonymous):

\[11x+510-17x=450\] \[-6x+510=450\] \[-6x=-60\] one more step and you are done

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

or 10 lbs

OpenStudy (anonymous):

ten, not minus ten ten pounds of the $11 coffee, the other 20 of the $17 coffee